### v1

# RAPTOR: Practical Numerical Profiling of Scientific Applications

Abstract.
The proliferation of low-precision units in modern high-performance architectures increasingly burdens domain scientists. Historically, the choice in HPC was easy: can we get away with 32 bit floating-point operations and lower bandwidth requirements, or is FP64 necessary? Driven by Artificial Intelligence, vendors introduced novel low-precision units for vector and tensor operations, and FP64 capabilities stagnate or are reduced. This is forcing scientists to re-evaluate their codes, but a trivial search-and-replace approach to go from FP64 to FP16 will not suffice.
We introduce RAPTOR: a numerical profiling tool to guide scientists in their search for code regions where precision lowering is feasible. Using LLVM, we transparently replace high-precision computations using low-precision units, or emulate a user-defined precision. RAPTOR is a novel, feature-rich approach—with focus on ease of use—to change, profile, and reason about numerical requirements and instabilities, which we demonstrate with four real-world multi-physics Flash-X applications.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Mathematics of computing Arbitrary-precision arithmeticccs: General and reference Empirical studiesccs: Computing methodologies Massively parallel and high-performance simulations
Figure 1. Qualitative visualization of bubble deformation and splitting over time ($t$) for the Bubble test case at Reynolds number $\text{Re}=3500$, performed using incompressible multiphase fluid dynamics solvers in Flash-X. The contours depict the zero level of the level-set function, $\phi=0$, which defines the air-water interface ($\phi>0$ in the air phase and $\phi<0$ in the water phase). Adaptive mesh refinement (AMR) levels $\Omega_{M}$, $\Omega_{M-1}$, and $\Omega_{M-2}$ are shown and labeled in the left panel. AMR dynamically refines the mesh near the interface to accurately capture interface dynamics. We use RAPTOR to truncate FP operations in the advection and diffusion modules of the Navier-Stokes solver following two different strategies. In the first strategy (Trunc. Everywhere), truncation is applied throughout the entire mesh. The second strategy selectively applies truncation in each mesh cell using a cut-off at the $M-l$ AMR level (where $l=1,2,\ldots$). Insets show zoomed-in views of the liquid-gas interface for simulations using 4 bit and 12 bit mantissas, highlighting the effect of reduced precision on the final simulation outcome. Further details in §6.2.
Table 1. Categorization and comparison of RAPTOR to existing approaches and tools. These categories are used: (A) Algorithmic mixed precision, (B) Automatically changing precision, (C) System software-enabled precision changes, (D) Application-granularity, (E) Wrapper and emulator, (F) Precision format, and (G) Observe behavior w/o changing precision. Feature details are as follows: Full app. runs indicates support for unmodified software w/o infeasible runtime increase [^6] Binary for tools which are able to instrument executables.. Dynamic truncation means that truncation of variables can be made algorithmically dependent on the simulation state77footnotemark: 7. Flexible formats stands for arbitrary precision support. Scoped truncation lets the user mark a function/region and the tool truncates the entire call stack below88footnotemark: 8. Granular truncation means that individual operations can be in-/excluded from truncation99footnotemark: 9. Error tracking traces errors through the application code1010footnotemark: 10. Non-$\nabla$ code means that arbitrary (even non-differentiable) code is supported.
Approach Category Feature set Ref.
Full app.
truncation
Dynamic
trunc.
Flexible
formats
Scoped
trunc.
Granular
trunc.
Error
tracking
Non-$\nabla$
Code
Supported
languages1111footnotemark: 11
[…]
1. Introduction
Scientific computing has traditionally relied upon double precision floating-point arithmetic to ensure numerical accuracy and stability across a wide range of problems. However, the growing availability of hardware optimized for lower precision—such as single or even half precision—presents an opportunity to significantly reduce energy consumption and improve computational efficiency. The core dilemma lies in balancing performance gains with the need for reliable scientific results. Some numerical methods are more amenable than others to systematic reduction of precision and can rely on a robust mathematical foundation for maintaining confidence in the outcome. Of these, numerical linear algebra has witnessed the highest level of activity. Several other areas such as many ordinary- and partial-differential equation (ODE and PDE) solvers do not have such mathematically founded support for lowering precision. While many parts of a simulation may tolerate lower precision without compromising accuracy, identifying these regions is nontrivial and highly problem-dependent. Using full precision everywhere guarantees numerical robustness but can be wasteful, whereas an indiscriminate use of reduced precision risks introducing errors that propagate and distort the final outcome as Figure 1 illustrates for a rising bubble simulation.
Here, the incompressible multiphase fluid-dynamics problem involves the deformation and splitting of an air bubble rising in a pool of water. Figure 1 visualizes the evolution of the liquid-gas interface and highlights how different mantissa truncations and numerical precisions (4 bit vs. 12 bit mantissa) affect interface dynamics. A zoomed-in view emphasizes these differences. A full analysis of the experiment is provided in Sections 4 to 6. This example demonstrates that for such workloads the challenge is to develop intelligent strategies and tools that enable selective precision tuning without undermining the fidelity of scientific simulations.
In this paper, we introduce a numerical profiling methodology and accompanying new tool, called RAPTOR, that can numerically profile code execution and enable scientists to infer the impact of reduced precision in an actionable way. We apply RAPTOR to a diverse set of application configurations in Flash-X (Dubey et al., 2022), a multiphysics multidomain application software, where mathematical guidance typically predicts that double precision is needed almost everywhere. We use scientific intuition to selectively reduce precision in code components and observe the impact through the profiles generated by RAPTOR. These experiments effectively demonstrate a methodology where generated insights enable meaningful cost-benefit analysis for running simulation campaigns where mathematics does not provide any guidance.
With focus on the scientific applications community, our paper makes the following contributions:
  • We introduce RAPTOR to easily profile and alter floating-point precision in user-specified code regions for C, C++, and/or Fortran; including MPI support. [Section 3]
  • We devise a methodology and demonstrate how RAPTOR can be used to experiment and reason about scientific application behavior to make informed choices about where to lower (or increase) the precision. [Section 4Section 6]
  • We discuss potential use cases of RAPTOR in characterizing precision needs of HPC workloads for hardware co-design in exploring the optimal distribution and the types of necessary floating-point units. [Section 7.2]
2. Background, Motivation, and Related Work
The exploration of mixed-precision arithmetic in scientific computing has gathered attention in recent years due to the potential to enhance computational performance and efficiency without compromising the accuracy of the results. By strategically combining different levels of precision within numerical algorithms, researchers aim to leverage the speed and reduced resource consumption of lower-precision computations while maintaining the numerical stability provided by higher-precision calculations.
Abdelfattah et al. (Abdelfattah et al., 2021) surveyed mixed-precision numerical linear algebra methods, discussing their theoretical foundations and practical applications. The study highlights how low-/mixed-precision techniques, e.g., iterative refinement, can achieve the accuracy of double-precision. Similarly, Higham et al. (Higham and Mary, 2022) provide an extensive review of mixed-precision algorithms in numerical linear algebra, covering a range of problems from factorization to solving linear systems. Bhola et al. (Bhola and Duraisamy, 2024) conducted a comprehensive study on rounding error analysis for mixed-precision arithmetics. Their work provides both deterministic and probabilistic approaches to quantify accumulated rounding errors in various operations.
While mixed-precision arithmetic has seen growing adoption in areas like linear algebra, machine learning, and some PDE solvers, several areas of numerical methods remain relatively underexplored because they involve complex numerical stability and error propagation behavior, making precision tuning less straightforward. Examples include methods such as multigrid (MG) and adaptive mesh refinement (AMR), nonlinear solvers and root-finding algorithms, general-purpose nonlinear solvers, spectral methods used in high-accuracy solutions of differential equations. These methods typically demand high-precision throughout and have not been a focus of mixed-precision research due to the assumption that double precision is required for accuracy preservation.
It is, however, possible to leverage scientific and physical intuition to predict if—and where—it might be feasible to lower the precision. Since there is no robust mathematical framework to validate such assumptions, there is a need for tools that enable interactive and responsive exploration of truncation strategy and provide detailed and immediate insights into the impact of precision changes. We have conducted a review of the existing tools, and define six categories of what and how the approach/tool tackles the numerical analysis and a set of desired features. Table 10 shows these categorizations and feature comparisons, but also highlights that none of the existing tools are feature-complete or easy to apply to an existing, potentially large code base and scientific workloads.
Only NEAT (Barati et al., 2021) offers a broad set of comparable features. However, it lacks the scoped precision changes which we rely on to analyze subsections of code or individual physics kernels in a multi-physics application. ADAPT (Menon et al., 2018) can be used to collect floating-point precision sensitivity profiles, but only works for differentiable applications. This downside is missing from CADNA (Jézéquel et al., 2021), FPSpy (Dinda et al., 2020), and FPVM(Dinda et al., 2022), but none of them support dynamic truncations (i.e., changes in mantissa length depending on the application’s state or other compile-time or runtime conditions). Both GPU-FPX (Li et al., 2023) and GPUMixer (Laguna et al., 2019) exhibit similar shortcomings to the other approaches, on top of being exclusively applicable to GPGPU codes. The remaining tools we are aware of, e.g., the ones developed by Jost et al. (Jost et al., 2021), Gu et al. (Gu et al., 2020), as well as Precimonious (Rubio-González et al., 2013) and Puppeteer (Parasyris et al., 2022), are unable to track numerical errors throughout the code (similar to our mem-mode, cf. Section 6.3) or correlate them to individual source locations. RAPTOR is designed to overcome these limitations.
3. RAPTOR: a Floating-Point Profiling Tool
Our goal for RAPTOR is to numerically profile applications or regions of interest, to determine whether the workload is amenable to execution with lowered precision throughout, under certain conditions only, or not all. We aim to make the iterative process of making changes to an application based on a hypothesis, profiling, and confirming the results as frictionless as possible.
3.1. Conceptual Design
Configurations
RAPTOR can scope its transformation to three levels: function, file, and program; and it has two modes of operation: op-mode and mem-mode. The supported configurations of scope and mode are shown in Table 2(a). The mem-mode requires more intervention from the user, and in turn, provides richer information about the computation (cf. Section 3.5 for further details).
Components
RAPTOR consists of two main components: a compiler instrumentation pass and a supporting runtime. The compiler pass inserts calls into the runtime at appropriate places in the user’s program. The runtime executes floating-point operations in the instructed precisions and collects data for analysis. Their place in the compilation pipeline is shown in Figure 2(a).
(a) RAPTOR’s components
Scope op-mode mem-mode
Function
Fully
Auto.
Semi
Auto.
File
Fully
Auto.
N/A
Program
Fully
Auto.
N/A
Features op-mode mem-mode
Profile Local Extensive
Precision
Increase
N/A Supported
(a) Supported scopes and features for each mode
Figure 2. Overview of RAPTOR and config. matrix
Figure 3. Usage of RAPTOR to go from original code to either op-mode (right) or mem-mode (left).
Figure 4. RAPTOR’s transformation pass replaces FP operations with runtime calls. We also implement an optimizations which allocates a temporary MPFR variable once to avoid re-allocation for each operation (cf. Figure 5a).
3.2. Usage
In general, there are flags [^12] The flags differ by compiler/linker and will be explained in the paper’s artifacts. which must be added to load our plugin into the compiler and adjust its optimization pipeline. The simplest way to use RAPTOR is at the file or program scope. This only requires an additional flag for the compilation command, e.g.: --raptor-truncate-all=64_to_5_14;32_to_3_8. The flag instructs RAPTOR to perform a truncation on the currently compiled files for the instructed truncations (in this case 64 bit width floating-point operations to a type with 5 bit exponent and 14 bit mantissa, and analogously for 32 bit width). If specified for all files, we achieve full-program truncation. To use function scope truncation, the code needs to be instrumented as shown in Figure 3. The user can request RAPTOR for a truncated version of a function by using the _raptor_trunc_func_[op,mem] functions, depending on the desired mode. The parameters specify the function and floating-point width to be truncated alongside the target exponent/mantissa lengths [^13] Currently, we require the exponent/mantissa to be compile-time constants.. The returned value is a function with the same type as the input function. Figure 3c shows how the mem-mode truncation can be used. Using mem-mode is more contrived and will be covered in detail in Section 3.5.
3.3. Instrumentation Pass
The RAPTOR pass is implemented as an instrumentation pass on the LLVM intermediate representation (or LLVM IR) (Lattner and Adve, 2004). LLVM is used by the clang (C/C++) and flang compilers among others. This allows us to handle all of Fortran, C, and C++ under a common representation suitable for compiler transformation and analysis [^14] We are providing examples in C++, however RAPTOR’s concepts apply equivalently to C and Fortran as well (with syntactical differences).. We build on top of Enzyme’s (Moses and Churavy, 2020) infrastructure and make use of its instrumentation and plugin utilities [^15] Enzyme is an auto-differentiation framework for LLVM..
Transformation
Figure 4a shows an example of the transformation when the user has requested a function scope truncation. It is shown in C++ for illustration purposes, however, the actual code being transformed is in LLVM IR. First, the pass finds all functions that the user requested be truncated and the truncation configuration (in this case, the user requested the function foo to be truncated). Then, we find all transitively called functions (in this case, foo and bar), and for each of them, we replace all operations on floating-point numbers with calls into our runtime. Here, by utilizing features available in LLVM, we can recognize floating-point arithmetic and functions in math libraries, such as the standard C and C++ math library. All affected functions are cloned transformations in order to preserve the behavior of unrelated code that uses them. When the requested scope is file- or program-level, our pass applies the same transformation to the floating-point operations to all functions, without the special handling required for function-scope truncation. This pass is implemented as an LLVM pass plugin, allowing it to be used in compilers such as clang or flang via their plugin interface in supported linkers (lld and gold), and as a standalone pass.
Preserving Compiler Optimizations
The stage of the optimization pipeline in which the pass runs dictates how closely the truncated operations match the optimized version of the floating-point operation. Compilers have various optimizations that affect floating-point computation, e.g., vectorization or fusion. However, they can only handle operations they are aware of. For example, in Figure 3a, the code contains floating-point arithmetic and a sqrt call to the libc math library, both of which are known to the optimizer, so it can treat them appropriately. However, once the code has been transformed (Figure 4a) into calls into our runtime, the optimizer no longer understands the semantics of the operation and no further floating-point operation optimization is possible.
Ideally, our pass would run after generic floating-point optimizations are done to preserve as many compiler optimizations as possible. However, due to current limitations (see Section 7.3), we are not able to handle vectorized instructions, which means we opt to run our pass before the vectorization passes.
Linking
If a call to a function with no definition is encountered while attempting to truncate, we are not able to instrument any operations inside it. For this reason, we configure the compiler(s) to use link-time optimizations (LTO), which merges all LLVM modules from different files into one file. We run our pass after the merge, which allows us to analyze the entire call graph [^16] Calls to pre-compiled external libraries will be ignored and we emit a warning., see Figure 2(a).
3.4. Runtime
The RAPTOR runtime executes floating-point operations in the specified precisions and collects data for analysis. This can be done using hardware, when the target precision is available on the CPU. Alternatively, for arbitrary precisions, we use GNU’s MPFR library (Fousse et al., 2007) for floating-point emulation, which provides feature rich and efficient implementation for arithmetic and mathematical functions. Figure 5 sketches the MPFR-based wrappers in RAPTOR.
The runtime also keeps track of how many floating-point operations are executed and how much memory is accessed in truncated and non-truncated regions. This information can be used to model an estimated speedup as shown in Section 7.2.
3.5. Operation Modes
As mentioned earlier, RAPTOR can operate in two modes: op-mode and mem-mode. The main difference is how the runtime behaves.
Op-Mode
Figure 5. Implementation of the RAPTOR runtime.
In this mode, the floating-point values that cross the boundary between the RAPTOR runtime and the user code (i.e. parameters to the runtime calls and return values) are represented using the pre-truncated floating-point type. For example, in the case in Figure 4a, the arguments to the _raptor_add_f32 function and its return value will be valid floating-point numbers of type float. Figure 5a shows the runtime implementation. Each time we perform any floating-point operation on these values, they first need to be transformed to the desired truncated precision (mpfr_set), the operations performed in truncated precision, and then expanded again to the original floating-point type (mpfr_get) [^17] This approach ensures correctness in adjunct code regions, e.g., calls to printf.. Since the values are converted back to the original type after each operation, there is no way of tracking how values flow through the computation. So, this mode is useful for collecting statistics about errors in individual operations (hence, op-mode), and the number thereof.
Mem-Mode
In this mode, values are not converted back to floats after each operation. Instead, the MPFR representation is memorized (hence, mem-mode) and is maintained between operations. Therefore, mem-mode allows precision increases, not only truncations.
Figure 5b shows a simplified version of the mem-mode runtime implementation. We use the bits in the floating-point type to store an integer identifier, which is used to recover a struct (_raptor_fp) containing the MPFR variable among other (customizable) bookkeeping data. Here, we add a double-precision shadow variable to the struct and update it with full-precision operations alongside the truncated MPFR variable. Hence, we can monitor the deviation between truncated variables and the (FP64) shadow variables for every operation, allowing us to set/monitor error threshold (cf. Section 6.3) and correlate them back to source code locations.
[…]
MPFR Optimization
[…]
4. Precision Considerations in Non-linear Multi-Physics Workloads
To evaluate RAPTOR in the upcoming sections, we require a set of realistic application use cases. Our intent is to select a class of applications and numerical methods that do not have mathematical foundations that could otherwise help us understand and predict behavior under mixed-precision. We use Flash-X (Dubey et al., 2022), a scientific software that is used in multiple science domains. Flash-X is particularly well suited for this study because its target applications are non-linear and coupled multi-physics problems which are sensitive to numerical perturbations. Additionally, Flash-X uses Adaptive Mesh Refinement (AMR), a computational technique used to dynamically adjust the resolution of a simulation mesh based on the evolving features of the solution. Instead of using a uniform grid everywhere, AMR refines the mesh in regions where fine detail is needed (e.g., near steep gradients or discontinuities such as shocks and interfaces) and coarsens it where the solution is smooth. This allows simulations to achieve high accuracy while reducing computational cost by concentrating resources where they matter most. In the context of this study, the use of AMR permits us to exercise scientific intuition for adjusting the precision because AMR relies heavily on user-defined or computed thresholds for the refinement.
(a) The radial shock in the Sedov blast simulation
(b) Shock along vertical plane is Sod shock-tube simulation
Figure 6. Compressible hydrodynamics workloads with varying AMR block sizes. Each block contains a fixed number of grid cells. The background color qualitatively represents pressure variations, with darker regions indicating higher pressure and lighter regions indicating lower pressure.
4.1. Experiment Design
Flash-X uses a version of AMR where the physical domain is divided into blocks that are spatially organized in an octree. The physical size of the blocks at a given level on the tree is identical; blocks a level above are twice the size along each dimension, while those at the level below are half the size. Figure 6(a) illustrates this for a 2D domain. The number of cells is identical in every block.
Our first set of experiments (see Section 6.1 and Section 6.2) is predicated on the intuition that the refinement criteria mentioned above ensure that all blocks that need the highest level of accuracy will be at the finest refinement level in the octree. The blocks at coarser levels in the tree will have smoother solutions implying that they may have a potential for adjusting the precision. In applications that simulate shock waves, the highest refinement will follow the progress of the shock, while in the applications which simulate multiphase flows, the highest resolution will follow the phase boundaries. We can explore this hypothesis by applying truncation on blocks that are at higher levels in the tree. We can investigate the impact of precision by varying the levels at which we commence the truncation, or by truncating specific physics solvers, and correlating the resulting error to our knowledge of the physical regimes being explored in each simulation. These experiments use RAPTOR’s op-mode with function and file-level scoping.
With another set of experiments (see Section 6.3) we will demonstrate the use of RAPTOR where no prior assumptions are made based on scientific intuition. Here, we operate in a numerical debugging mode akin to performance debugging. We use mem-mode with function and file-level scoping for these experiments. In this mode, RAPTOR flags the operations which deviate from a reference value by more than a predefined threshold. Working backwards from the flagged operations one can roll back to full precision recursively until the desired accuracy is restored.
4.2. Flash-X Workloads and Hypotheses
Flash-X can be configured for different applications with different underlying physics and solvers. Here, we select applications in two different physical regimes, i.e., with compressible and incompressible hydrodynamics solvers. In workloads with compressible hydrodynamics, shocks (sharp, nearly discontinuous change in the properties of fluid over a very small region of space) develop due to discontinuities in the initial conditions. We select two applications with different shock profiles. The first case, Sedov blast wave (Sedov, 1993), is initialized with a pressure spike in the center of the domain where the shock moves out in the radial direction. The regions away from the shock are more or less quiescent (see Figure 6(a)). The second case is the Sod shock tube (Sod, 1978) where there is a jump in density along the vertical plane. The shock wave moves in one direction while the rarefaction wave moves in the opposite direction, see Figure 6(b). For these workloads, we theorize:
Hypothesis 1: In Sedov, some regions close to the shock have a higher resolution than the physical conditions strictly demand. Thus, we expect that reducing precision in all but the most refined blocks would not significantly impact the quality of the results. In Sod, the solution profile is less sharp and stretches across coarser blocks, and hence we expect reduced precision to have a larger impact on the results.
Another study in the compressible regime is the cellular detonation (Timmes et al., 2000) (Cellular) with nuclear burning and an equation of state (EOS) suitable for stellar interiors. The domain is initialized with pure carbon which is perturbed to ignite the nuclear fuel, producing an over-driven detonation that propagates along the $x$-axis. The EOS module uses a table of Helmholtz free energy with discrete values, and extrapolates them to match the conditions in the domain. For Cellular, we intend to explore the possibility of using lower precision in a solver other than hydro in a multiphysics scenario:
Hypothesis 2: EOS, used in the simulation, is table-based, and is therefore the most likely candidate for reducing precision.
The application in the incompressible regime is a rising bubble benchmark (Bubble). This solver employs a fractional-step projection method to evolve the velocity field and a sharp-interface ghost fluid method to model multiphase interfacial dynamics (Dhruv, 2024, 2021). Here, a circular air bubble of diameter $d=1.0$ is initialized with its center at the origin, $(0,0)$, within a two-dimensional rectangular domain. The air–water interface is tracked using a level-set function, $\phi$, where $\phi>0$ denotes the air phase, $\phi<0$ corresponds to the water phase, and $\phi=0$ defines the air-water interface.
This Bubble simulation is governed by key dimensionless parameters for fluid and flow configuration. With $\rho^{\prime}=1000$ and $\mu^{\prime}=100$, we denote the density and viscosity ratios between water and air, respectively. These two parameters control the scaling of the fluid forces between the two phases. Furthermore, the Reynolds number (Re), Froude number (Fr), and Weber number (We) characterize the relative importance of inertial, gravitational, and surface tension forces. These numbers are set for water phase and scaled in air by the density and viscosity ratios, and we use $\text{Fr}=1$ and $\text{We}=125$. This configuration corresponds to a benchmark case reported in (Dhruv, 2024, 2023).
We use the solution at $t=3$ computed at $\text{Re}=35$ as the starting point for a series of simulations in which truncation is applied to both advection and diffusion operators of the Navier-Stokes solver. The advection terms are discretized using a fifth-order Weighted Essentially Non-Oscillatory (WENO) scheme, while a second-order central difference scheme is used for diffusion. These simulations are performed at $\text{Re}=3500$. The choice of higher Re for truncation tests was motivated by the need to accelerate bubble splitting and deformation in a shorter time period from $t=3$ to $t=4$. Our working hypothesis for this Bubble test is as follows:
Hypothesis 3: The required precision will align with the AMR refinement strategy, specifically that reducing precision in lower-resolution blocks will not significantly degrade the overall quality of the simulation. We also expect that numerical precision will influence the evolution of the bubble interface, affecting deformation, splitting, and shape over time. Since the precision in velocity calculations directly impacts bubble dynamics, we expect noticeable differences as the precision varies.
5. Experimental Setup
Here, we briefly describe details about the sub-components of RAPTOR, the execution environment, and how to read the figures.
(a) Between the 1st and 2nd panels, the error drops by seven orders of magnitude for mantissas smaller than 18 bits and remains constant at $10^{-15}$ for mantissas larger than 18 bits. Restricting truncation to coarser levels in 3rd/4th panels does not significantly change the error plot. Gray lines shadow other panels’ error to ease the comparison.
(b) Between the 1st and 2nd panels, the error drops slightly for mantissas smaller than 18 bits and matches the baseline elsewhere. In the 3rd panel, the error improves by half an order of magnitude for mantissas smaller than 48 bits. The sharp error drop for small mantissas is caused by the AMR algorithm refining all blocks.
Figure 7. Truncating hydrodynamics in the Sedov blast wave simulation (left) and Sod shock tube simulation (right). The primary $y$-axis shows the $L1$ error norm of the fluid density compared to the full-precision simulation, as computed by Flash-X’ serial output comparison utility sfocu. Results are plotted with lines (not points) for clarity. Gray lines, showing the errors from the other panels, are included to emphasize small differences in error from one panel to the next. Black vertical lines mark half, float, and double precision mantissa sizes with 10, 23, and 52 bits, respectively. In these plots, each panel represents a different refinement cutoff level beyond which truncation was disabled (i.e., $M-0$: truncate everything; $M-1$: disable truncation for most refined AMR level; $M-2$: disable truncation for two most refined levels; etc.). The bars in the background of each plot show the number of truncated operations (red) and full-precision operations (orange) stacked on top. As the refinement cutoff level is coarsened, the portion of truncated operations shrinks as truncation is restricted to coarser levels. Section 7.2 demonstrates how the operation counts can be used with a simple performance model to predict speedups. For small mantissas towards the left of each panel, the operation counts tend to deviate from the mean. The counts are perturbed when the AMR algorithm refines the grid to compensate for inaccuracies caused by aggressive truncation.
5.1. Versioning and Execution Environment
For LLVM, we use tag 20.1.0-init and enhance it with a patch to add function wrappers for Fortran. Our fork of Enzyme is based on the trunc-trace branch. We utilize Spack (Gamblin et al., 2015) to install MPFR in version 4.2.1 and to install a Bubble dependency, called Hypre (v2.31.0). We compile additional Flash-X dependencies (HDF5 v1.14.6, OpenMPI v5.0.6, AmReX v24.08, MA28 v1.0.0) needed for the experiments. To gain access to the compressible hydrodynamics applications, we fork Flash-X’ main branch, state of 2025/01/24202020Flash-X carries an Apache 2.0 license but is maintained privately; Please email flash-x@lists.cels.anl.gov with your Github ID to get access (reference: #ec2a750).. For the Bubble application, we fork Flash-X’ state of 2025/04/14212121For reference: #cf6c018..
We run all experiments in the default configuration for Flash-X: CPU-only and MPI-only. The Sod, Sedov, and Cellular applications execute sufficiently fast, and hence we use a single MPI process. The Bubble workload is computationally more demanding, requiring us to execute it with 32 MPI ranks, while still confining it to a single compute node. However, the parallelization across ranks does not affect the outcome of our experiments [^22] No MPI collectives are called within truncated sections; the domain is split over MPI processes and the physics routines that we truncate operate locally in each cell.. Our cluster’s compute nodes are equipped with dual-socket AMD EPYC 7773X processors and 1024 GiB of DDR4 main memory.
We employ the same methodology for each experiment: First, we evenly lower precision in a given module across the entire mesh to set the baseline of expected behavior. Then, we start with a very small mantissa and gradually increase its size getting an error estimate for each instance. Once the baseline is established we repeat the above process by lowering precision on blocks in the target physics solver at levels $M-1$, $M-2$, and $M-3$ respectively where $M$ is the maximum refinement level.
6. RAPTOR Case Studies: Experimental Results and Insights for the four AMR Workloads
Hereafter, we describe our results from analyzing the Flash-X workloads with RAPTOR. We explore three truncation modes: (1) Global truncation, where numerical values are truncated uniformly across the entire domain, (2) Selective truncation with AMR, where truncation is applied only on levels coarser than ${M-l}$, with $l$ controlling the cutoff depth in the AMR hierarchy, and (3) Selective truncation of a complete physics module.
6.1. Compressible hydrodynamics applications
Figures 7(a) and 7(b) show the results of the hydro experiments. As expected, the Sedov problem exhibits a very robust tolerance to lowered precision at $M-1$ as demonstrated in the second panel of Figure 7(a), which shows a much lower error for any size of mantissa compared to the error shown in the first panel. For mantissas with less than 28 bits, excluding the finest AMR blocks from truncation results in an error reduction of more than 7 orders-of-magnitude. For mantissas larger than 28 bits, the error remains constant and approaches the error measured for full truncation. This error floor appears as the simulation result drifts over time due to small inaccuracies in truncated floating-point operations. Excluding the second level of AMR blocks from truncation does not significantly improve the error of the simulation, as demonstrated by the unchanging error in Panel 3. Only when we exclude the finest three levels of AMR blocks from truncation do we see that the constant error floor drops by another half order of magnitude as shown in Panel 4.
The colored bars in the background of the figure panels count the number of truncated and full-precision floating-point operations occurring for each simulation. In the fourth panel, the workload is comprised almost entirely of full-precision operations, with truncated operations making up less than $0.4\%$. In the third and second panel, the truncated operations make up progressively larger portions of the workload as the Hydro module is truncated for more blocks in the domain. Finally, in the first panel, the truncated operations make up more than 80% of the floating-point workload. Additionally, for mantissas smaller than 10 bits, the bars show an instability, with the overall flop count going up and down. This occurs because the low-precision operations are influencing the decisions of the AMR algorithm, which is responsible for dynamically refining the grid where needed. To be clear, it is not the algorithm itself which is working with truncated precision. Rather, the algorithm notices imprecise blocks and decides to refine them in order to maintain numerical accuracy and convergence in the simulation.
Figure 7(b) shows the error plot for the Sod problem. Comparing the first two panels, we observe that excluding the finest AMR blocks from truncation results in an improvement in the overall error for mantissa sizes smaller than 18 bits. This improvement results in an almost order-of-magnitude difference in error between the first and second panel for simulations using 4 bits of mantissa. Panel 3 shows that excluding another level of AMR blocks from truncation results in an error improvement across the board, with the difference starting at about half an order-of-magnitude for large mantissas with 48 bits and increasing to almost an order-of-magnitude for small mantissas with 10 bits.
Excluding refined blocks from truncation improves the error more for the Sedov problem than it does for the Sod problem, validating Hypothesis 1. For Sedov, running just the finest blocks at full precision results in an improvement of 7 orders-of-magnitude compared to truncating all blocks. For Sod, the same strategy decreases the error by an order-of-magnitude at most. The figure for Sod has one less panel, because eventually no leaf blocks remain at the $M-3$ level in the simulation and therefore there is no error.
In Figure 7(b), an anomalous behavior is visible for mantissas bits between 4-6, where the errors are comparable to those obtained for mantissas with 20 bits. The reason for that anomaly is written in the logfile of the Flash-X run, i.e., it shows a higher number of leaf blocks (the blocks on which the solution evolves) which implies that many more blocks meet the refinement criterion at such low mantissa bits. Consequently, any advantage from lowering the precision is more than offset by the increase of refined blocks. It highlights the importance of a careful cost-benefit analysis of any performance-related tweaks in the application and showcases how RAPTOR can assist with such an analysis.
With Cellular we turn our attention to the possibility of truncating an entire physics module. The EOS module is the obvious choice here because the ordinary differential equations (ODEs) in the Burn module are particularly stiff and sensitive to numerical perturbation. Hypothesis 2 is based on the assumption that the computation can work with reduced precision because it extrapolates from a table look-up. However, this hypothesis is proven to be incorrect because the Newton-Raphson algorithm, used to extrapolate from the table look-up values to actual conditions in the simulation, does not converge within the specified number of iterations when the mantissa is truncated to less than 42 bits. In an attempt to get around this problem, we decrease the tolerance for convergence and increase the permitted number of iterations. Yet, we fail to get convergence for any meaningful workload, providing us with a counterexample for our hypothesis. This experiment in particular highlights the usefulness of RAPTOR in numerical profiling because scientific intuition is not always reliable. Any direct modification of code to run such experiments would be time consuming and not as informative.
6.2. Incompressible multiphase flow application
For this experiment, we focus on the behavior of the bubble dynamics under varying levels of precision. We omit the error curves because they do not provide any additional insights compared to the results already shown for Sedov and Sod in Figure 7. Figure 1 provides snapshots of the Rising Bubble benchmark at $\text{Re}=3500$, showcasing the deformation and splitting of an air bubble rising through quiescent water. The interface is captured using a level-set formulation, with $\phi=0$ denoting the liquid-gas boundary. The AMR hierarchy is centered around the interface to capture these fine-scale features with high fidelity. The key aim of this experiment is to test Hypothesis 3, by evaluating how different truncation strategies and numerical precision levels for the advection and diffusion operators affect the evolution of the bubble interface.
We compare results obtained with low (4 bit) and moderate (12 bit) precision. The zoomed-in insets in Figure 1 highlight the qualitative differences in interface topology under these settings. With aggressive truncation and low precision, the interface develops visible artifacts (see the $t=3.5$ and $t=4$ panel) particularly related to the shape of parent and satellite bubbles during and after break-up where the solution is highly sensitive to numerical inaccuracies. In contrast, with moderate precision and more selective truncation we are able to preserve the shape and physical accuracy of the interface evolution without requiring FP64 computations.
6.3. Mem-mode debugging
In this section, we demonstrate how the debugging output generated by RAPTOR in mem-mode can identify sections of code that may be unstable after truncatation. A domain scientist can use this information to understand which parts of the code contribute most to the overall accuracies when truncated.
To recap Section 3.5: mem-mode uses shadow variables that track an reference result for every operation as if the entire application had been run in FP64 up to that point. This feature can be used to pinpoint exactly where calculations start to deviate from the reference (above a threshold). RAPTOR flags these occurrences of inexact operation results, groups them by (debug) location in the code, and dumps the collected statistics when instructed by the user. In practice, certain instances of operations get flagged more often than others and flagged operations tend to be close together in the code. Thus, the output of mem-mode builds a heatmap of code locations that do not react well to truncation.
We apply mem-mode debugging to the Sedov problem with a newer, more modular hydrodynamics solver, Spark (Couch et al., 2021). The modularity of Spark makes it easier to ascribe a certain task to each part of the code. For example, two important components of the solver are the reconstruction algorithm and the Riemann solver. The reconstruction algorithm approximates the variation of the solution within each cell with a profile instead of using a constant value. The Riemann solver handles discontinuous solutions in shocks. For this experiment, we keep the timestep of the solver constant to ensure that the dynamic time-stepping algorithm does not interfere with our experiment by compensating for inaccuracies resulting from truncation.
Table 2. Numerically debugging Sedov with mem-mode
$L1$ error norm
Excluded modules density x-velocity Truncated FP ops
Baseline 8.113e-4 4.338e-3 90.6%
[…]
Recon, Riemann $\blacktriangle$ 9.300e-4 $\blacktriangle$ 6.929e-3 9.1%
[…]
We start by truncating the entirety of the hydrodynamics module and compute the resulting errors for density and x-velocity using the sfocu utility, which is designed to verify the correctness of outputs from Flash-X simulations against reference benchmarks (public known good solutions). These errors serve as a baseline for the rest of the experiment, as summarized in Table 2. The goal is to improve the error by selectively fencing off sections of the truncated module and running them at full double precision.
After we truncate the hydrodynamics, RAPTOR flags a number of operations in the reconstruction algorithm. This can be interpreted in two ways. Either, (a) the reconstruction algorithm does not respond well to truncation and should therefore be run at full precision, or (b) another part of the code is responsible for introducing small floating-point errors that only get flagged when they are amplified in the reconstruction algorithm, for example due to multiplication or cancellation. If we exclude reconstruction from truncation (but still truncate the remaining hydrodynamics code), then the errors for both density and x-velocity decrease by a small amount. Now, a number of new operations get flagged in the Riemann solver and a single operation gets flagged in the function responsible for updating the solution of the simulation. Adding the Riemann solver to the list of excluded functions drastically worsens errors, while adding the update function leaves the errors essentially unchanged.
These results indicate that no specific part of the Spark solver is responsible for numerical sensitivity in the Sedov workload and it is difficult to decide in advance whether a code section should be truncated or not. Instead, an effective truncation strategy will dynamically turn on and off based on the local smoothness of the solution in the physical regime. This small example reaffirms the need for a tool, such as RAPTOR, that can help the users to understand and reason about the parts of a code that may contribute the most to floating-point errors.
7. Discussion and Outlook
We used RAPTOR to test our hypotheses about how to truncate a number of non-linear multi-physics workloads while keeping errors in check. For Sedov, we found that restricting truncation to coarse blocks worked well to keep errors low. Meanwhile for Sod, the same strategy resulted in much higher errors. These two results match Hypothesis 1. However, our Hypothesis 2 was falsified when we tried to truncate the EOS module in Cellular and found that it only converges when we run it with double precision. Finally, we evaluated the effect of precision on the evolution of the Bubble workload and found that position and shape of the split bubbles depends on the strategy and level of truncation, as the scientific intuition anticipated.
7.1. Overhead
Table 3. Slowdown of RAPTOR in practice
Runtime (s) Overhead (in $\times$)
Truncated FP ops naive opt. naive opt.
Sedov in op-mode (12 bit)
M-0 86.3% 4781 1883 91.9 36.3
M-1 31.0% 1734 672 33.3 13.0
M-2 13.6% 794 323 15.3 6.2
M-3 0.37% 75 59 1.4 1.1
Sedov in op-mode with operation counting (12 bit)
M-0 86.3% 5229 2336 24.2 45.0
M-1 31.0% 2088 1005 9.7 19.4
M-2 13.6% 1075 590 5.0 11.4
M-3 0.4% 285 279 1.3 5.4
Sedov in mem-mode (12 bit)
Truncate Hydro 90.6% 556 N/A 148 N/A
Exclude Recon [^23] Exclusion in mem mode is handled dynamically in the runtime, so both entries have comparable overhead. 17.8% 543 N/A 147 N/A
RAPTOR’s overhead for typical executions is shown in Table 3. The overhead roughly correlates to the proportion of operations we truncate, and in the simplest op-mode it can get up to 36$\times$ slower (with lower overheads for lower proportions). Since our tool is not meant to be used in production runs, and is only meant for experimentation and reasoning about precision, we believe that the overhead is within acceptable ranges that make it useful.
Figure 8. Estimated speedup of Sod (cf. §4.2 and Figure 7(b)) for different truncation strategies, according to our hardware model under a compute-bound or memory-bound assumption. Here, our model predicts a compute-bound scenario.
7.2. Use in Hardware Co-design
We want to also highlight other adjunct usecases for RAPTOR, e.g., as part of the hardware co-design (w.r.t. floating-point capabilities).
Floating-Point Unit (FPU) Model
Assume, we have a simple model for a hypothetical processor with a fixed portion of chip area dedicated to floating-point processing. Because the workloads, which we investigated earlier, are all double precision floating-point, and we truncate a portion of them to a lower precision, we can assume that our CPU contains FPUs for double precision and one lower precision. For simplicity, we will assume that the areas dedicated to each unit remain the same.
Table 4. Performance density of FPUs for various precisions (data from FPNew (Mach et al., 2020))
FP Type GFLOP/s Area (kGE) Perf. density (normalized)
fp64 (11, 52) 3.17 53 1.00
fp32 (8, 23) 6.33 40 2.65
fp16 (5, 10) 12.67 29 7.30
fp8 (5, 2) 25.33 23 18.41
We estimate how much processing power in floating-point operations per second (FLOP/s) a unit area of chip footprint provides for a given precision [^24] For simplicity, we assume that floating point units with different precision are completely independent. In reality, floating point units for different precisions often share some components.. We get an estimate for the performance density of an FPU using data from an open-source RISCV FPU implementation (i.e., FPNew (Mach et al., 2020)), which is summarized in Table 4. We extrapolate these values to get a performance density estimate for FPUs of any given precision.
Our theoretical CPU has a separate FPU for the two supported precisions (called $M_{\mathrm{dbl}}$ and $M_{\mathrm{low}}$), which, respectively, occupy an area of $A_{\mathrm{dbl}}$ and $A_{\mathrm{low}}$ and have a performance density of $P_{\mathrm{dbl}}$ and $P_{\mathrm{low}}$ operations per second per unit area (obtained above using extrapolation). Therefore, the processing capability for a specific precision is bound by a peak of $A_{i}\times P_{i}$ operations per unit time. To estimate how much area each of the two FPUs will occupy, we assume a typical CPU configuration where we have double and single precision capabilities with a ratio of $1:2$ (e.g. Fugaku’s A64FX (fug, [n. d.])). Using the estimated performance densities $P_{i}$, and the ratio of the compute power, we get the ratio for the areas $A_{\mathrm{dbl}}:A_{\mathrm{low}}=1.39$.
Next, we consider how to evaluate the time spent for floating-point operations of a specific computation which contains a number of floating-point operations $N_{i}$ for each precision $i\in M$. We assume that there is no parallelism across the FPUs and only a single type of FPU is in use at any given time, which matches the way we truncate all operations in specific regions in this work.
Consequently, the time spent for the computation in a precision $i$ is $\frac{N_{i}}{A_{i}\times P_{i}}$ and the time spent for the entire computation is given by the following: $\sum_{i\in\mathrm{dbl},\mathrm{low}}\frac{N_{i}}{A_{i}\times P_{i}}$. We use this formula to obtain an estimated speedup for a computation in case it is compute-bound.
Memory Model
As mentioned in Section 3.4, we collect information about memory accesses that happen in truncated and non-truncated regions. We use these numbers to get a rough estimate of how much memory movement can be saved when floating-point values are truncated to a lower precision. In case of memory-bound code, we assume that the runtime is a linear function of the required memory movement and obtain an estimated speedup.
Roofline Model
To get a prediction for whether a computation is compute- or memory-bound, we build a roofline model of our processor, assuming a bandwidth of $1024$ GB/s (e.g., Fugaku’s (fug, [n. d.])).
Results
Using the above model and the data collected from our experiments described earlier, we obtain estimated speedups for different configurations of Sod and present the results in Figure 8. Due to this being a computation-heavy application that has a high operation intensity (operations per byte), our roofline model predicted a compute-bound scenario. The irregularities around the small mantissas are due to the AMR behavior changing and resulting in a different ratios of truncated and non-truncated operations. With insight from domain scientists on acceptable levels of error obtained from Figure 7(b), this information can be used to obtain an optimal FPU configuration for hardware co-design. Given that RAPTOR simplifies the collection of such information, collaborating with scientists for gathering data on the numerical behavior of software can become a powerful way to enable supercomputing centers to make informed decisions about future procurements appropriate for their workloads.
7.3. Current Limitations
For our proof-of-concept, we omitted a few quality-of-life features. For example, not all elementary functions are implemented, but adding additional functions is trivial if MPFR already supports them, otherwise minor engineering is required. RAPTOR does not truncate function calls to external, pre-compiled libraries [^25] This can be mitigated by making the function definitions available, either by providing source code or static libraries compiled with link-time optimization (LTO). or calls to dynamic function pointers. We have not tested OpenMP support or thread-safety in general, original floating-point operands have to be FP32 or FP64, and we do not support handwritten vector assembly. These limitations can be overcome with small engineering effort. As shown in Figure 4, the user has to manually change a few lines of code. To improve the user experience, RAPTOR can be enhanced for easier region annotation (e.g. #pragma trunc...) and to support function filtering using a configuration file (similar to profilers). Programming languages other than C/C++ and Fortran could be supported if they lower to LLVM IR. Lastly, our truncation parameters are compile-time constants, but deciding the truncation level at runtime can be achieved by compiling multiple function pointers for different truncations and conditionally using them.
8. Conclusion
This paper introduces a novel numerical profiling approach and accompanying tool, called RAPTOR, which overcomes the limitations of state-of-the-art tools in this category. Using an LLVM compiler pass and MPFR, we are able to transparently change the floating-point precision by modifying the exponent and/or mantissa of selected code regions. We create hypotheses for multiple, different workloads in the Flash-X framework for multi-physics simulations. Using our RAPTOR tool, we demonstrate how we are able to analyze and confirm/falsify theses hypotheses based on collected data for the non-linear solvers which are usually hard to reason about. Our hypotheses for Sod, Sedov and Bubble experiments prove to be correct. However, our intuition turns out to be incorrect for the Cellular experiment. Currently, we can apply our tool to C/C++ and Fortran codes, and hence it supports the majority of (large-scale) scientific workloads run on supercomputers. While we outlined a few limitations, none of them are fundamental to our approach, and, as we mention, these can be mitigated with additional engineering effort. RAPTOR is a step forward towards reasoning about mixed/low precision in our simulations, and we hope that it finds wide adoption in the scientific-computing community. We make our RAPTOR tool accessible at https://github.com/RAPTOR-FP/RAPTOR for others to use.

### v2

@@
 Historically, the choice in HPC was easy: can we get away with 32 bit floating-point operations and lower bandwidth requirements, or is FP64 necessary?
-Driven by Artificial Intelligence, vendors introduced novel low-precision units for vector and tensor operations, and FP64 capabilities stagnate or are reduced.
-This is forcing scientists to re-evaluate their codes, but a trivial search-and-replace approach to go from FP64 to FP16 will not suffice.
+Driven by Artificial Intelligence, vendors introduce novel low-precision units for vector and tensor operations, and FP64 capabilities stagnate or are reduced.
+This forces scientists to re-evaluate their codes, but a trivial search-and-replace approach to go from FP64 to FP16 will not suffice.
 We introduce RAPTOR: a numerical profiling tool to guide scientists in their search for code regions where precision lowering is feasible.
@@
 RAPTOR is a novel, feature-rich approach—with focus on ease of use—to change, profile, and reason about numerical requirements and instabilities, which we demonstrate with four real-world multi-physics Flash-X applications.
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Mathematics of computing Arbitrary-precision arithmeticccs: General and reference Empirical studiesccs: Computing methodologies Massively parallel and high-performance simulations
+Mixed precision, low precision, numerical profiling, error tracking, simulation accuracy, multiphysics, LLVM, MPFR
+journalyear: 2025copyright: ccconference: The International Conference for High Performance Computing, Networking, Storage and Analysis; November 16–21, 2025; St Louis, MO, USAbooktitle: The International Conference for High Performance Computing, Networking, Storage and Analysis (SC ’25), November 16–21, 2025, St Louis, MO, USAdoi: 10.1145/3712285.3759810isbn: 979-8-4007-1466-5/2025/11ccs: Mathematics of computing Arbitrary-precision arithmeticccs: General and reference Empirical studiesccs: Computing methodologies Massively parallel and high-performance simulations
 Figure 1.
@@
 Everywhere), truncation is applied throughout the entire mesh.
-The second strategy selectively applies truncation in each mesh cell using a cut-off at the $M-l$ AMR level (where $l=1,2,\ldots$).
+The second strategy selectively applies truncation in each mesh cell using a cut-off at the MlM-l AMR level (where $l=1,2,\ldots$).
 Insets show zoomed-in views of the liquid-gas interface for simulations using 4 bit and 12 bit mantissas, highlighting the effect of reduced precision on the final simulation outcome.
-Further details in §6.2.
+Further details in §6.2.:
 Table 1.
@@
 Feature details are as follows: Full app.
-runs indicates support for unmodified software w/o infeasible runtime increase [^6] Binary for tools which are able to instrument executables..
-Dynamic truncation means that truncation of variables can be made algorithmically dependent on the simulation state77footnotemark: 7.
+truncation indicates support for unmodified software w/o infeasible runtime increase.
+[^2] E.g., Flash-X can not realistically be analyzed with ADAPT without applying kernel extraction or similar techniques to isolate individual regions of interest in the workload.
+Dynamic truncation means that truncation of variables can be made algorithmically dependent on the simulation state.
+[^3] E.g., NEAT has functionality to change behavior based on the calling context.
 Flexible formats stands for arbitrary precision support.
-Scoped truncation lets the user mark a function/region and the tool truncates the entire call stack below88footnotemark: 8.
-Granular truncation means that individual operations can be in-/excluded from truncation99footnotemark: 9.
-Error tracking traces errors through the application code1010footnotemark: 10.
+Scoped truncation lets the user mark a function/region and the tool truncates the entire call stack below.
+[^4] E.g., tools making use of LLVM IR have this functionality.
+Granular truncation means that individual operations can be in-/excluded from truncation.
+[^5] E.g., the approach by Gu et al.
+considered individual flops for tuning.
+Error tracking traces errors through the application code.
+[^6] E.g., GPUMixer demonstrated this by using shadow variables.
 Non-$\nabla$ Code means that arbitrary (even non-differentiable) code is supported.
@@
 Dynamic
-trunc.
+truncation
 Flexible
@@
 Scoped
-trunc.
+truncation
 Granular
-trunc.
+truncation
 Error
@@
 Supported
-languages1111footnotemark: 11
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+languages [^7] Binary for tools which are able to instrument executables.
+ADAPT B , , Σ , Σ Σ , C, C++, Fortran (Menon et al., 2018)
+CADNA C Σ , , N/A Σ Σ Σ Ada, C, Fortran (Jézéquel et al., 2021)
+FPSpy G Σ , , N/A N/A , Σ Binary (Dinda et al., 2020)
+FPVM D Σ , Σ , , Σ Σ Binary (Dinda et al., 2022)
+GPU-FPX G Σ , N/A N/A , Σ Σ GPU Binary (Li et al., 2023)
+GPUMixer B , , , , , Σ , CUDA (Laguna et al., 2019)
+Jost et al. C, E, F Σ , Σ N/A Σ , Σ C (Jost et al., 2021)
+Gu et al. E , , , , Σ , Σ C, C++ (Gu et al., 2020)
+NEAT E Σ Σ Σ , Σ , Σ Binary (Barati et al., 2021)
+Precimonious A , N/A , Σ , , Σ C (Rubio-González et al., 2013)
+Puppeteer D Σ Σ N/A , Σ , , C, C++ (Parasyris et al., 2022)
+RAPTOR B, C, E Σ Σ Σ Σ Σ Σ Σ C, C++, Fortran
 1. Introduction
-Scientific computing has traditionally relied upon double precision floating-point arithmetic to ensure numerical accuracy and stability across a wide range of problems.
+Scientific computing has traditionally relied on double precision floating-point arithmetic to ensure numerical accuracy and stability across a wide range of problems.
 However, the growing availability of hardware optimized for lower precision—such as single or even half precision—presents an opportunity to significantly reduce energy consumption and improve computational efficiency.
@@
 While many parts of a simulation may tolerate lower precision without compromising accuracy, identifying these regions is nontrivial and highly problem-dependent.
-Using full precision everywhere guarantees numerical robustness but can be wasteful, whereas an indiscriminate use of reduced precision risks introducing errors that propagate and distort the final outcome as Figure 1 illustrates for a rising bubble simulation.
+Using full precision everywhere guarantees numerical robustness but can be wasteful, whereas an indiscriminate use of reduced precision risks introducing errors that propagate and distort the final outcome.
+Figure 1 illustrates this for a rising bubble simulation.
 Here, the incompressible multiphase fluid-dynamics problem involves the deformation and splitting of an air bubble rising in a pool of water.
@@
 12 bit mantissa) affect interface dynamics.
-A zoomed-in view emphasizes these differences.
+Zoomed-in views emphasize these differences.
 A full analysis of the experiment is provided in Sections 4 to 6.
@@
 We apply RAPTOR to a diverse set of application configurations in Flash-X (Dubey et al., 2022), a multiphysics multidomain application software, where mathematical guidance typically predicts that double precision is needed almost everywhere.
-We use scientific intuition to selectively reduce precision in code components and observe the impact through the profiles generated by RAPTOR.
+We use scientific intuition to selectively reduce precision in code components and observe the impact through the profiles generated with the help of RAPTOR.
 These experiments effectively demonstrate a methodology where generated insights enable meaningful cost-benefit analysis for running simulation campaigns where mathematics does not provide any guidance.
 With focus on the scientific applications community, our paper makes the following contributions:
-  • We introduce RAPTOR to easily profile and alter floating-point precision in user-specified code regions for C, C++, and/or Fortran; including MPI support. [Section 3]
+  • We introduce RAPTOR [^1] https://github.com/RIKEN-RCCS/RAPTOR to easily profile and alter floating-point precision in user-specified code regions for C, C++, and Fortran; including GPU and OpenMP support.
+[Section 3]
   • We devise a methodology and demonstrate how RAPTOR can be used to experiment and reason about scientific application behavior to make informed choices about where to lower (or increase) the precision.
@@
 Examples include methods such as multigrid (MG) and adaptive mesh refinement (AMR), nonlinear solvers and root-finding algorithms, general-purpose nonlinear solvers, spectral methods used in high-accuracy solutions of differential equations.
-These methods typically demand high-precision throughout and have not been a focus of mixed-precision research due to the assumption that double precision is required for accuracy preservation.
+These methods typically demand high precision throughout and have not been a focus of mixed-precision research due to the assumption that double precision is required for accuracy preservation.
 It is, however, possible to leverage scientific and physical intuition to predict if—and where—it might be feasible to lower the precision.
-Since there is no robust mathematical framework to validate such assumptions, there is a need for tools that enable interactive and responsive exploration of truncation strategy and provide detailed and immediate insights into the impact of precision changes.
+Since there is no robust mathematical framework to validate such assumptions, there is a need for tools that enable interactive and responsive exploration of truncation strategies and provide detailed and immediate insights into the impact of precision changes.
 We have conducted a review of the existing tools, and define six categories of what and how the approach/tool tackles the numerical analysis and a set of desired features.
@@
 3. RAPTOR: a Floating-Point Profiling Tool
-Our goal for RAPTOR is to numerically profile applications or regions of interest, to determine whether the workload is amenable to execution with lowered precision throughout, under certain conditions only, or not all.
+Our goal for RAPTOR is to numerically profile applications or regions of interest to determine whether the workload is amenable to execution with lowered precision throughout, under certain conditions only, or not all.
 We aim to make the iterative process of making changes to an application based on a hypothesis, profiling, and confirming the results as frictionless as possible.
@@
 RAPTORcan scope its transformation to three levels: function, file, and program; and it has two modes of operation: op-mode and mem-mode.
-The supported configurations of scope and mode are shown in Table 2(a).
+The supported configurations of scope and mode are shown in Figure 2(b).
 The mem-mode requires more intervention from the user, and in turn, provides richer information about the computation (cf.
@@
 RAPTORconsists of two main components: a compiler instrumentation pass and a supporting runtime.
-The compiler pass inserts calls into the runtime at appropriate places in the user’s program.
+The compiler pass inserts calls to the runtime at appropriate places in the user’s program.
 The runtime executes floating-point operations in the instructed precisions and collects data for analysis.
-Their place in the compilation pipeline is shown in Figure 2(a).
+The placement of the components in the compilation pipeline is shown in Figure 2(a).
 (a) RAPTOR’s components
@@
 N/A Supported
-(a) Supported scopes and features for each mode
-Figure 2. Overview of RAPTOR and config. matrix
-Figure 3. Usage of RAPTOR to go from original code to either op-mode (right) or mem-mode (left).
+(b) Supported scopes and features for each mode
+Figure 2. Overview of RAPTOR and config. matrix:
+Figure 3. Usage of RAPTOR to go from original code to either op-mode (right) or mem-mode (left).:
+{subcaptiongroup}
+[…]
+[…]
 Figure 4.
 RAPTOR’s transformation pass replaces FP operations with runtime calls.
-We also implement an optimizations which allocates a temporary MPFR variable once to avoid re-allocation for each operation (cf.
-Figure 5a).
+We also implement an optimization which allocates a temporary MPFR variable once to avoid re-allocation for each operation (cf.
+Figure 5).
+:
+{subcaptiongroup}
+[…]
+[…]
+Figure 5. Implementation of the RAPTOR runtime.:
 3.2. Usage
-In general, there are flags [^12] The flags differ by compiler/linker and will be explained in the paper’s artifacts.
+In general, there are flags [^8] The flags differ by compiler/linker and are detailed in the paper’s artifacts.
 which must be added to load our plugin into the compiler and adjust its optimization pipeline.
@@
 To use function scope truncation, the code needs to be instrumented as shown in Figure 3.
-The user can request RAPTOR for a truncated version of a function by using the _raptor_trunc_func_[op,mem] functions, depending on the desired mode.
-The parameters specify the function and floating-point width to be truncated alongside the target exponent/mantissa lengths [^13] Currently, we require the exponent/mantissa to be compile-time constants..
-The returned value is a function with the same type as the input function.
-Figure 3c shows how the mem-mode truncation can be used.
-Using mem-mode is more contrived and will be covered in detail in Section 3.5.
+The user can request a truncated version of a function using the _raptor_trunc_func_[op,mem] functions, depending on the desired mode.
+The parameters specify the function and floating-point width to be truncated alongside the target exponent/mantissa lengths.
+[^9] Currently, we require the exponent/mantissa to be compile-time constants.
+The return value is a function with the same type as the input function.
+Using mem-mode truncation, shown in Figure 3c, is more contrived and will be covered in detail in Section 3.5.
 3.3. Instrumentation Pass
-The RAPTOR pass is implemented as an instrumentation pass on the LLVM intermediate representation (or LLVM IR) (Lattner and Adve, 2004).
+The RAPTOR pass is implemented as an instrumentation pass on the LLVM intermediate representation (LLVM IR) (Lattner and Adve, 2004).
 LLVM is used by the clang (C/C++) and flang compilers among others.
-This allows us to handle all of Fortran, C, and C++ under a common representation suitable for compiler transformation and analysis [^14] We are providing examples in C++, however RAPTOR’s concepts apply equivalently to C and Fortran as well (with syntactical differences)..
-We build on top of Enzyme’s (Moses and Churavy, 2020) infrastructure and make use of its instrumentation and plugin utilities [^15] Enzyme is an auto-differentiation framework for LLVM..
+This allows us to handle all of Fortran, C, and C++ under a common representation suitable for compiler transformation and analysis.
+[^10] We are providing examples in C++, however RAPTOR’s concepts apply equivalently to C and Fortran as well (with syntactical differences).
+We build on top of Enzyme’s (Moses and Churavy, 2020) infrastructure [^11] Enzyme is an auto-differentiation framework for LLVM.
+and make use of its instrumentation and plugin utilities.
 Transformation
-Figure 4a shows an example of the transformation when the user has requested a function scope truncation.
-It is shown in C++ for illustration purposes, however, the actual code being transformed is in LLVM IR.
+Figure 4 shows an example of the transformation when the user has requested a function scope truncation.
+It is shown in C++ for illustrative purposes, but the actual code being transformed is in LLVM IR.
 First, the pass finds all functions that the user requested be truncated and the truncation configuration (in this case, the user requested the function foo to be truncated).
-Then, we find all transitively called functions (in this case, foo and bar), and for each of them, we replace all operations on floating-point numbers with calls into our runtime.
-Here, by utilizing features available in LLVM, we can recognize floating-point arithmetic and functions in math libraries, such as the standard C and C++ math library.
+Then, the pass finds all transitively called functions (in this case, foo and bar), and for each of them, replaces all operations on floating-point numbers with calls into the RAPTOR runtime.
+By utilizing features available in LLVM, we can recognize floating-point arithmetic and functions in math libraries, such as the standard C and C++ math library.
 All affected functions are cloned transformations in order to preserve the behavior of unrelated code that uses them.
-When the requested scope is file- or program-level, our pass applies the same transformation to the floating-point operations to all functions, without the special handling required for function-scope truncation.
-This pass is implemented as an LLVM pass plugin, allowing it to be used in compilers such as clang or flang via their plugin interface in supported linkers (lld and gold), and as a standalone pass.
+When the requested scope is file or program-level, our pass applies the same transformation to the floating-point operations of all functions, without the special handling required for function-scope truncation.
+This is implemented as an LLVM pass plugin, which can be used in compilers such as clang or flang via their plugin interface in supported linkers (lld and gold), and as a standalone pass.
 Preserving Compiler Optimizations
@@
 For example, in Figure 3a, the code contains floating-point arithmetic and a sqrt call to the libc math library, both of which are known to the optimizer, so it can treat them appropriately.
-However, once the code has been transformed (Figure 4a) into calls into our runtime, the optimizer no longer understands the semantics of the operation and no further floating-point operation optimization is possible.
-Ideally, our pass would run after generic floating-point optimizations are done to preserve as many compiler optimizations as possible.
+However, once the code has been transformed to call into our runtime (Figure 4), the optimizer no longer understands the semantics of the operation and no further floating-point operation optimization is possible.
+Ideally, our pass would run after all generic floating-point optimizations are done to preserve as many compiler optimizations as possible.
 However, due to current limitations (see Section 7.3), we are not able to handle vectorized instructions, which means we opt to run our pass before the vectorization passes.
@@
 For this reason, we configure the compiler(s) to use link-time optimizations (LTO), which merges all LLVM modules from different files into one file.
-We run our pass after the merge, which allows us to analyze the entire call graph [^16] Calls to pre-compiled external libraries will be ignored and we emit a warning., see Figure 2(a).
+We run our pass after the merge, which allows us to analyze the entire call graph, [^12] Calls to pre-compiled external libraries are ignored and RAPTOR emits a warning.
+see Figure 2(a).
 3.4. Runtime
@@
 This can be done using hardware, when the target precision is available on the CPU.
-Alternatively, for arbitrary precisions, we use GNU’s MPFR library (Fousse et al., 2007) for floating-point emulation, which provides feature rich and efficient implementation for arithmetic and mathematical functions.
+Alternatively, for arbitrary precisions, we use the GNU MPFR library (Fousse et al., 2007) for floating-point emulation, which provides feature rich and efficient implementations of arithmetic and mathematical functions.
 Figure 5 sketches the MPFR-based wrappers in RAPTOR.
@@
 Op-Mode
-Figure 5. Implementation of the RAPTOR runtime.
 In this mode, the floating-point values that cross the boundary between the RAPTOR runtime and the user code (i.e.
 parameters to the runtime calls and return values) are represented using the pre-truncated floating-point type.
-For example, in the case in Figure 4a, the arguments to the _raptor_add_f32 function and its return value will be valid floating-point numbers of type float.
-Figure 5a shows the runtime implementation.
-Each time we perform any floating-point operation on these values, they first need to be transformed to the desired truncated precision (mpfr_set), the operations performed in truncated precision, and then expanded again to the original floating-point type (mpfr_get) [^17] This approach ensures correctness in adjunct code regions, e.g., calls to printf..
+For example, in the case in Figure 4, the arguments to the _raptor_add_f32 function and its return value will be valid floating-point numbers of type float.
+Figure 5 shows the runtime implementation.
+Each time we perform any floating-point operation on these values, they first need to be transformed to the desired truncated precision (mpfr_set), the operations performed in truncated precision, and then expanded again to the original floating-point type (mpfr_get).
+[^13] This approach ensures correctness in adjunct code regions, e.g., calls to printf.
 Since the values are converted back to the original type after each operation, there is no way of tracking how values flow through the computation.
 So, this mode is useful for collecting statistics about errors in individual operations (hence, op-mode), and the number thereof.
+[…]
 Mem-Mode
@@
 Therefore, mem-mode allows precision increases, not only truncations.
-Figure 5b shows a simplified version of the mem-mode runtime implementation.
+Figure 5 shows a simplified version of the mem-mode runtime implementation.
 We use the bits in the floating-point type to store an integer identifier, which is used to recover a struct (_raptor_fp) containing the MPFR variable among other (customizable) bookkeeping data.
 Here, we add a double-precision shadow variable to the struct and update it with full-precision operations alongside the truncated MPFR variable.
-Hence, we can monitor the deviation between truncated variables and the (FP64) shadow variables for every operation, allowing us to set/monitor error threshold (cf.
+Hence, we can monitor the deviation between truncated variables and the (FP64) shadow variables for every operation, allowing us to set/monitor error thresholds (cf.
 Section 6.3) and correlate them back to source code locations.
 […]
-MPFR Optimization
-[…]
+(a) The radial shock in the Sedov blast simulation
+(b) Shock along vertical plane in the Sod shock-tube simulation
+Figure 6.
+Compressible hydrodynamics workloads with varying AMR block sizes.
+Each block contains a fixed number of mesh cells.
+The background color qualitatively represents pressure variations, with darker regions indicating higher pressure and lighter regions indicating lower pressure.:
+3.6. Compatibility
+RAPTORrecognizes OpenMP directives and correctly truncates operations within nested OpenMP parallel constructs within the same device.
+In addition to truncating applications for CPU, RAPTOR can truncate code for GPUs (AMD and NVidia).
+We have successfully tested OpenMP offloading and CUDA code.
+Due to the lack of GPU support in arbitrary precision libraries (such as MPFR), truncation is currently only possible to native types available in the GPU hardware, such as float or half precision floating-point.
+In general, RAPTOR’s op-mode and MPI do not interfere with one another and truncation continues to work for any application with one or more MPI ranks.
+Most MPI operations only involve message passing and therefore require no special handling.
+However, RAPTOR does not implicitly truncate MPI reductions or any other MPI functions that perform floating-point operations.
+If e.g.
+truncated MPI Allreduce is needed, a custom reduction operation can be implemented, which in turn can be truncated using RAPTOR.
+Mem-mode can only be used on shared-memory systems and without MPI reductions.
+[^16] In op-mode, reductions do not get truncated automatically.
+In mem-mode, reductions will crash the application when they change the pointers to the shadow storage.
+Otherwise, additional MPI communication and handling would be required to transfer the shadow storage and adjust the pointers.
 4. Precision Considerations in Non-linear Multi-Physics Workloads
 To evaluate RAPTOR in the upcoming sections, we require a set of realistic application use cases.
-Our intent is to select a class of applications and numerical methods that do not have mathematical foundations that could otherwise help us understand and predict behavior under mixed-precision.
+Our intent is to select a class of applications and numerical methods that do not have mathematical foundations that could otherwise predict mixed-precision behavior.
 We use Flash-X (Dubey et al., 2022), a scientific software that is used in multiple science domains.
-Flash-X is particularly well suited for this study because its target applications are non-linear and coupled multi-physics problems which are sensitive to numerical perturbations.
+Flash-X is well suited for this study because its target applications are non-linear and coupled multi-physics problems which are sensitive to numerical perturbations.
 Additionally, Flash-X uses Adaptive Mesh Refinement (AMR), a computational technique used to dynamically adjust the resolution of a simulation mesh based on the evolving features of the solution.
@@
 This allows simulations to achieve high accuracy while reducing computational cost by concentrating resources where they matter most.
-In the context of this study, the use of AMR permits us to exercise scientific intuition for adjusting the precision because AMR relies heavily on user-defined or computed thresholds for the refinement.
-(a) The radial shock in the Sedov blast simulation
-(b) Shock along vertical plane is Sod shock-tube simulation
-Figure 6.
-Compressible hydrodynamics workloads with varying AMR block sizes.
-Each block contains a fixed number of grid cells.
-The background color qualitatively represents pressure variations, with darker regions indicating higher pressure and lighter regions indicating lower pressure.
+Here, the use of AMR permits us to exercise scientific intuition for adjusting the precision because AMR relies heavily on user-defined or computed thresholds for the refinement.
 4.1. Experiment Design
 Flash-Xuses a version of AMR where the physical domain is divided into blocks that are spatially organized in an octree.
-The physical size of the blocks at a given level on the tree is identical; blocks a level above are twice the size along each dimension, while those at the level below are half the size.
+The physical size of the blocks at a given level on the tree is identical; blocks one level above are twice the size along each dimension, while those one level below are half the size.
 Figure 6(a) illustrates this for a 2D domain.
@@
 In applications that simulate shock waves, the highest refinement will follow the progress of the shock, while in the applications which simulate multiphase flows, the highest resolution will follow the phase boundaries.
-We can explore this hypothesis by applying truncation on blocks that are at higher levels in the tree.
 We can investigate the impact of precision by varying the levels at which we commence the truncation, or by truncating specific physics solvers, and correlating the resulting error to our knowledge of the physical regimes being explored in each simulation.
@@
 We select two applications with different shock profiles.
-The first case, Sedov blast wave (Sedov, 1993), is initialized with a pressure spike in the center of the domain where the shock moves out in the radial direction.
+The first case, Sedov blast wave (Sedov, 2018), is initialized with a pressure spike in the center of the domain where the shock moves out in the radial direction.
 The regions away from the shock are more or less quiescent (see Figure 6(a)).
@@
 For Cellular, we intend to explore the possibility of using lower precision in a solver other than hydro in a multiphysics scenario:
-Hypothesis 2: EOS, used in the simulation, is table-based, and is therefore the most likely candidate for reducing precision.
+Hypothesis 2: The EOS, used in the simulation, is table-based and is therefore the most likely candidate for reducing precision.
 The application in the incompressible regime is a rising bubble benchmark (Bubble).
@@
 The air–water interface is tracked using a level-set function, $\phi$, where $\phi>0$ denotes the air phase, $\phi<0$ corresponds to the water phase, and $\phi=0$ defines the air-water interface.
-This Bubble simulation is governed by key dimensionless parameters for fluid and flow configuration.
+The Bubble simulation is governed by key dimensionless parameters for fluid and flow configuration.
 With $\rho^{\prime}=1000$ and $\mu^{\prime}=100$, we denote the density and viscosity ratios between water and air, respectively.
@@
 These numbers are set for water phase and scaled in air by the density and viscosity ratios, and we use $\text{Fr}=1$ and $\text{We}=125$.
-This configuration corresponds to a benchmark case reported in (Dhruv, 2024, 2023).
+This configuration corresponds to a benchmark case reported in (Dhruv, 2024; Akash Dhruv, 2023).
 We use the solution at $t=3$ computed at $\text{Re}=35$ as the starting point for a series of simulations in which truncation is applied to both advection and diffusion operators of the Navier-Stokes solver.
@@
 5. Experimental Setup
-Here, we briefly describe details about the sub-components of RAPTOR, the execution environment, and how to read the figures.
+Here, we briefly discuss versions and details of the sub-components of RAPTOR. We also describe the execution environment and lay out our strategy for running the experiments.
 (a) Between the 1st and 2nd panels, the error drops by seven orders of magnitude for mantissas smaller than 18 bits and remains constant at $10^{-15}$ for mantissas larger than 18 bits.
-Restricting truncation to coarser levels in 3rd/4th panels does not significantly change the error plot.
-Gray lines shadow other panels’ error to ease the comparison.
+Restricting truncation to coarser levels in the 3rd and 4th panels does not significantly change the error plot.
 (b) Between the 1st and 2nd panels, the error drops slightly for mantissas smaller than 18 bits and matches the baseline elsewhere.
@@
 Black vertical lines mark half, float, and double precision mantissa sizes with 10, 23, and 52 bits, respectively.
-In these plots, each panel represents a different refinement cutoff level beyond which truncation was disabled (i.e., $M-0$: truncate everything; $M-1$: disable truncation for most refined AMR level; $M-2$: disable truncation for two most refined levels; etc.).
+In these plots, each panel represents a different refinement cutoff level beyond which truncation was disabled (i.e., M0M-0: truncate everything; M1M-1: disable truncation for most refined AMR level; M2M-2: disable truncation for two most refined levels; etc.).
 The bars in the background of each plot show the number of truncated operations (red) and full-precision operations (orange) stacked on top.
@@
 For small mantissas towards the left of each panel, the operation counts tend to deviate from the mean.
-The counts are perturbed when the AMR algorithm refines the grid to compensate for inaccuracies caused by aggressive truncation.
-5.1. Versioning and Execution Environment
-For LLVM, we use tag 20.1.0-init and enhance it with a patch to add function wrappers for Fortran.
+The counts are perturbed when the AMR algorithm refines the mesh to compensate for inaccuracies caused by aggressive truncation.
+:
+For LLVM, we use version 20, which offers a major improvement of the flang compiler over previous releases.
 Our fork of Enzyme is based on the trunc-trace branch.
-We utilize Spack (Gamblin et al., 2015) to install MPFR in version 4.2.1 and to install a Bubble dependency, called Hypre (v2.31.0).
+We utilize Spack (Gamblin et al., 2015) to install MPFR version 4.2.1 and to install a Bubble dependency, called Hypre (v2.31.0).
 We compile additional Flash-X dependencies (HDF5 v1.14.6, OpenMPI v5.0.6, AmReX v24.08, MA28 v1.0.0) needed for the experiments.
-To gain access to the compressible hydrodynamics applications, we fork Flash-X’ main branch, state of 2025/01/24202020Flash-X carries an Apache 2.0 license but is maintained privately; Please email flash-x@lists.cels.anl.gov with your Github ID to get access (reference: #ec2a750)..
-For the Bubble application, we fork Flash-X’ state of 2025/04/14212121For reference: #cf6c018..
+To gain access to the compressible hydrodynamics applications, we fork Flash-X’ main branch, state of 2025/01/24.171717Flash-Xcarries an Apache 2.0 license but is maintained privately; Please email flash-x@lists.cels.anl.gov with your Github ID to get access (reference: #ec2a750).
+For the Bubble application, we fork Flash-X’ state of 2025/04/14.181818For reference: #cf6c018.
 We run all experiments in the default configuration for Flash-X: CPU-only and MPI-only.
@@
 The Bubble workload is computationally more demanding, requiring us to execute it with 32 MPI ranks, while still confining it to a single compute node.
-However, the parallelization across ranks does not affect the outcome of our experiments [^22] No MPI collectives are called within truncated sections; the domain is split over MPI processes and the physics routines that we truncate operate locally in each cell..
+However, the parallelization across ranks does not affect the outcome of our experiments.
+[^19] No MPI collectives are called within truncated sections; the domain is split over MPI processes and the physics routines that we truncate operate locally in each cell.
 Our cluster’s compute nodes are equipped with dual-socket AMD EPYC 7773X processors and 1024 GiB of DDR4 main memory.
@@
 Then, we start with a very small mantissa and gradually increase its size getting an error estimate for each instance.
-Once the baseline is established we repeat the above process by lowering precision on blocks in the target physics solver at levels $M-1$, $M-2$, and $M-3$ respectively where $M$ is the maximum refinement level.
+Once the baseline is established, we repeat the above process by lowering precision on blocks in the target physics solver at levels M1M-1, M2M-2, and M3M-3 respectively where $M$ is the maximum refinement level.
 6. RAPTORCase Studies: Experimental Results and Insights for the four AMR Workloads
 Hereafter, we describe our results from analyzing the Flash-X workloads with RAPTOR.
-We explore three truncation modes: (1) Global truncation, where numerical values are truncated uniformly across the entire domain, (2) Selective truncation with AMR, where truncation is applied only on levels coarser than ${M-l}$, with $l$ controlling the cutoff depth in the AMR hierarchy, and (3) Selective truncation of a complete physics module.
+We explore three truncation modes: (1) Global truncation, where numerical values are truncated uniformly across the entire domain, (2) Selective truncation with AMR, where truncation is applied only on levels coarser than Ml{M-l}, with $l$ controlling the cutoff depth in the AMR hierarchy, and (3) Selective truncation of a complete physics module.
 6.1. Compressible hydrodynamics applications
 Figures 7(a) and 7(b) show the results of the hydro experiments.
-As expected, the Sedov problem exhibits a very robust tolerance to lowered precision at $M-1$ as demonstrated in the second panel of Figure 7(a), which shows a much lower error for any size of mantissa compared to the error shown in the first panel.
+As expected, the Sedov problem exhibits a very robust tolerance to lowered precision at M1M-1 as demonstrated in the second panel of Figure 7(a), which shows a much lower error for any size of mantissa compared to the error shown in the first panel.
 For mantissas with less than 28 bits, excluding the finest AMR blocks from truncation results in an error reduction of more than 7 orders-of-magnitude.
@@
 Finally, in the first panel, the truncated operations make up more than 80% of the floating-point workload.
-Additionally, for mantissas smaller than 10 bits, the bars show an instability, with the overall flop count going up and down.
-This occurs because the low-precision operations are influencing the decisions of the AMR algorithm, which is responsible for dynamically refining the grid where needed.
+Additionally, for mantissas smaller than 10 bits, the bars show irregularities, with the overall flop counts fluctuating.
+This occurs because the low-precision operations are influencing the decisions of the AMR algorithm, which is responsible for dynamically refining the mesh where needed.
 To be clear, it is not the algorithm itself which is working with truncated precision.
@@
 This improvement results in an almost order-of-magnitude difference in error between the first and second panel for simulations using 4 bits of mantissa.
-Panel 3 shows that excluding another level of AMR blocks from truncation results in an error improvement across the board, with the difference starting at about half an order-of-magnitude for large mantissas with 48 bits and increasing to almost an order-of-magnitude for small mantissas with 10 bits.
+Panel 3 shows that excluding another level of AMR blocks from truncation results in an error improvement across the board, with the difference starting at about half an order-of-magnitude for large mantissas with 48 bits and increasing to almost an order-of-magnitude for small mantissas with 10 bits or fewer.
 Excluding refined blocks from truncation improves the error more for the Sedov problem than it does for the Sod problem, validating Hypothesis 1.
@@
 For Sod, the same strategy decreases the error by an order-of-magnitude at most.
-The figure for Sod has one less panel, because eventually no leaf blocks remain at the $M-3$ level in the simulation and therefore there is no error.
-In Figure 7(b), an anomalous behavior is visible for mantissas bits between 4-6, where the errors are comparable to those obtained for mantissas with 20 bits.
-The reason for that anomaly is written in the logfile of the Flash-X run, i.e., it shows a higher number of leaf blocks (the blocks on which the solution evolves) which implies that many more blocks meet the refinement criterion at such low mantissa bits.
-Consequently, any advantage from lowering the precision is more than offset by the increase of refined blocks.
+The figure for Sod has one less panel, because eventually no leaf blocks remain at the M3M-3 level in the simulation, so no truncation occurs and therefore there is no error.
+In Figure 7(b), an anomalous behavior is visible for mantissas with 4-6 bits, where the errors are comparable to those obtained for mantissas with 20 bits.
+The reason for that anomaly is written in the logfile of the Flash-X run, i.e., it shows a higher number of leaf blocks (the blocks on which the solution evolves) which implies that many more blocks meet the refinement criterion with so few mantissa bits.
+Consequently, any advantage from lowering the precision further is more than offset by the increased number of refined blocks.
 It highlights the importance of a careful cost-benefit analysis of any performance-related tweaks in the application and showcases how RAPTOR can assist with such an analysis.
@@
 This experiment in particular highlights the usefulness of RAPTOR in numerical profiling because scientific intuition is not always reliable.
-Any direct modification of code to run such experiments would be time consuming and not as informative.
+Any direct modification of code to run such experiments with reduced precision would be time consuming and not as informative.
 6.2. Incompressible multiphase flow application
@@
 6.3. Mem-mode debugging
-In this section, we demonstrate how the debugging output generated by RAPTOR in mem-mode can identify sections of code that may be unstable after truncatation.
-A domain scientist can use this information to understand which parts of the code contribute most to the overall accuracies when truncated.
-To recap Section 3.5: mem-mode uses shadow variables that track an reference result for every operation as if the entire application had been run in FP64 up to that point.
+In this section, we demonstrate how the debugging output generated by RAPTOR in mem-mode can identify sections of code that may be unstable after truncation.
+A domain scientist can use this information to understand which parts of the code contribute most to the overall inaccuracies when truncated.
+To recap Section 3.5: mem-mode uses shadow variables that track a reference result for every operation as if the entire application had been run in FP64 up to that point.
 This feature can be used to pinpoint exactly where calculations start to deviate from the reference (above a threshold).
@@
 The Riemann solver handles discontinuous solutions in shocks.
-For this experiment, we keep the timestep of the solver constant to ensure that the dynamic time-stepping algorithm does not interfere with our experiment by compensating for inaccuracies resulting from truncation.
+For this experiment, we keep the timestep of the solver constant to ensure that the dynamic time-stepping algorithm does not compensate for inaccuracies resulting from truncation and thus skews our results.
 Table 2. Numerically debugging Sedov with mem-mode
@@
 If we exclude reconstruction from truncation (but still truncate the remaining hydrodynamics code), then the errors for both density and x-velocity decrease by a small amount.
-Now, a number of new operations get flagged in the Riemann solver and a single operation gets flagged in the function responsible for updating the solution of the simulation.
+Next, a number of new operations get flagged in the Riemann solver and a single operation gets flagged in the function responsible for updating the solution of the simulation.
 Adding the Riemann solver to the list of excluded functions drastically worsens errors, while adding the update function leaves the errors essentially unchanged.
@@
 Instead, an effective truncation strategy will dynamically turn on and off based on the local smoothness of the solution in the physical regime.
-This small example reaffirms the need for a tool, such as RAPTOR, that can help the users to understand and reason about the parts of a code that may contribute the most to floating-point errors.
+This small example reaffirms the need for a tool, such as RAPTOR, that can help users to understand and reason about the parts of a code that may contribute the most to floating-point errors.
 7. Discussion and Outlook
@@
 However, our Hypothesis 2 was falsified when we tried to truncate the EOS module in Cellular and found that it only converges when we run it with double precision.
-Finally, we evaluated the effect of precision on the evolution of the Bubble workload and found that position and shape of the split bubbles depends on the strategy and level of truncation, as the scientific intuition anticipated.
+Finally, we evaluated the effect of precision on the evolution of the Bubble workload and found that position and shape of the split bubbles depends on the strategy and level of truncation, as the scientific intuition anticipated in our Hypothesis 3.
 7.1. Overhead
@@
 Exclude Recon [^20] Exclusion in mem mode is handled dynamically in the runtime, so both entries have comparable overhead. 17.8% 543 N/A 147 N/A
-RAPTOR’s overhead for typical executions is shown in Table 3.
-The overhead roughly correlates to the proportion of operations we truncate, and in the simplest op-mode it can get up to 36$\times$ slower (with lower overheads for lower proportions).
-Since our tool is not meant to be used in production runs, and is only meant for experimentation and reasoning about precision, we believe that the overhead is within acceptable ranges that make it useful.
-Figure 8.
-Estimated speedup of Sod (cf.
-§4.2 and Figure 7(b)) for different truncation strategies, according to our hardware model under a compute-bound or memory-bound assumption.
-Here, our model predicts a compute-bound scenario.
+[…]
 7.2. Use in Hardware Co-design
-We want to also highlight other adjunct usecases for RAPTOR, e.g., as part of the hardware co-design (w.r.t. floating-point capabilities).
+We want to also highlight other adjunct usecases for RAPTOR, e.g., as part of hardware co-design (w.r.t. floating-point capabilities).
 Floating-Point Unit (FPU) Model
-Assume, we have a simple model for a hypothetical processor with a fixed portion of chip area dedicated to floating-point processing.
-Because the workloads, which we investigated earlier, are all double precision floating-point, and we truncate a portion of them to a lower precision, we can assume that our CPU contains FPUs for double precision and one lower precision.
-For simplicity, we will assume that the areas dedicated to each unit remain the same.
-Table 4. Performance density of FPUs for various precisions (data from FPNew (Mach et al., 2020))
+We assume a simple model for a hypothetical processor with a fixed portion of chip area dedicated to floating-point processing.
+Given that our investigated workloads are all double precision with some portion truncated to a lower precision, we can assume that our CPU contains FPUs for double precision and one lower precision.
+For simplicity, we assume that the areas dedicated to each unit remain the same.
+Table 4. Performance density of FPUs for various precisions (data from FPNew (Mach et al., 2021))
 FP Type GFLOP/s Area (kGE) Perf. density (normalized)
@@
 fp8 (5, 2) 25.33 23 18.41
-We estimate how much processing power in floating-point operations per second (FLOP/s) a unit area of chip footprint provides for a given precision [^24] For simplicity, we assume that floating point units with different precision are completely independent.
-In reality, floating point units for different precisions often share some components..
-We get an estimate for the performance density of an FPU using data from an open-source RISCV FPU implementation (i.e., FPNew (Mach et al., 2020)), which is summarized in Table 4.
+We estimate how much processing power in floating-point operations per second (FLOP/s) a unit area of chip footprint provides for a given precision [^21] For simplicity, we assume that floating point units with different precisions are completely independent.
+In reality, the units often share some components..
+We get an estimate for the performance density of an FPU using data from an open-source RISCV FPU implementation (i.e., FPNew (Mach et al., 2021)), which is summarized in Table 4.
 We extrapolate these values to get a performance density estimate for FPUs of any given precision.
@@
 Therefore, the processing capability for a specific precision is bound by a peak of $A_{i}\times P_{i}$ operations per unit time.
-To estimate how much area each of the two FPUs will occupy, we assume a typical CPU configuration where we have double and single precision capabilities with a ratio of $1:2$ (e.g.
-Fugaku’s A64FX (fug, [n.
+To estimate the area that each of the two FPUs will occupy, we assume a typical CPU configuration with double and single precision capabilities in a ratio of $1:2$ (e.g.
+Fugaku’s A64FX (Fujitsu, [n.
 d.])).
@@
 Roofline Model
-To get a prediction for whether a computation is compute- or memory-bound, we build a roofline model of our processor, assuming a bandwidth of $1024$ GB/s (e.g., Fugaku’s (fug, [n. d.])).
+To get a prediction for whether a computation is compute- or memory-bound, we build a roofline model of our processor, assuming a bandwidth of $1024$ GB/s (e.g., Fugaku’s (Fujitsu, [n. d.])).
+Figure 8.
+Estimated speedup of Sod (cf.
+§4.2 and Figure 7(b)) for different truncation strategies, according to our hardware model in compute-bound and memory-bound scenarios.
+Results are plotted with lines (not points) for clarity.
+Black vertical lines mark half, float, and double precision mantissa sizes with 10, 23, and 52 bits, respectively.
+Irregularities around the small mantissas are due to AMR changing the ratios of truncated and non-truncated operations.:
 Results
-Using the above model and the data collected from our experiments described earlier, we obtain estimated speedups for different configurations of Sod and present the results in Figure 8.
-Due to this being a computation-heavy application that has a high operation intensity (operations per byte), our roofline model predicted a compute-bound scenario.
-The irregularities around the small mantissas are due to the AMR behavior changing and resulting in a different ratios of truncated and non-truncated operations.
+Using the above model and the data collected from our experiments described in Section 6.1, we estimate speedups for different configurations of Sod and present the results in Figure 8.
+This is a computation-heavy application with a high operational intensity, so our roofline model predicts a compute-bound scenario.
+Predicted speedups for full truncation are $3.7\times$ for half precision and $2.2\times$ for single precision ($2.2\times$, $1.6\times$ memory-bound).
+The predicted speedups for M-1 and M-2 are lower than for M-0 because a smaller portion of the code is truncated.
+Consequently, speedups from low-precision operations impact the overall performance less.
+Irregularities around the small mantissas are caused by the adaptive behavior of AMR, which changes the ratio of truncated and non-truncated operations.
+This is reflected in the bar plots of Figure 7(b).
+For M-1, extra operations caused by AMR outweigh the speedup of the low-precision operations, resulting in net slowdowns for 4 and 5 bit mantissas.
 With insight from domain scientists on acceptable levels of error obtained from Figure 7(b), this information can be used to obtain an optimal FPU configuration for hardware co-design.
@@
 For our proof-of-concept, we omitted a few quality-of-life features.
-For example, not all elementary functions are implemented, but adding additional functions is trivial if MPFR already supports them, otherwise minor engineering is required.
+For example, not all elementary functions are implemented, but adding additional functions is trivial if MPFR already supports them.
 RAPTOR does not truncate function calls to external, pre-compiled libraries [^22] This can be mitigated by making the function definitions available, either by providing source code or static libraries compiled with link-time optimization (LTO).
 or calls to dynamic function pointers.
-We have not tested OpenMP support or thread-safety in general, original floating-point operands have to be FP32 or FP64, and we do not support handwritten vector assembly.
-These limitations can be overcome with small engineering effort.
-As shown in Figure 4, the user has to manually change a few lines of code.
+We have not tested thread-safety in general, original floating-point operands have to be FP16, FP32, or FP64, and we do not support handwritten vector assembly.
+Limitations outside of supporting handwritten assembly can be overcome with small engineering effort.
+As shown in Figure 3, the user has to manually change a few lines of code to use scoped truncation.
 To improve the user experience, RAPTOR can be enhanced for easier region annotation (e.g.
@@
 This paper introduces a novel numerical profiling approach and accompanying tool, called RAPTOR, which overcomes the limitations of state-of-the-art tools in this category.
-Using an LLVM compiler pass and MPFR, we are able to transparently change the floating-point precision by modifying the exponent and/or mantissa of selected code regions.
+Using an LLVM compiler pass and MPFR, we are able to transparently change the precision of floating-point operations in selected code regions.
 We create hypotheses for multiple, different workloads in the Flash-X framework for multi-physics simulations.
-Using our RAPTOR tool, we demonstrate how we are able to analyze and confirm/falsify theses hypotheses based on collected data for the non-linear solvers which are usually hard to reason about.
+Using our RAPTOR tool, we demonstrate how we are able to analyze and confirm/falsify these hypotheses based on collected data for the non-linear solvers which are usually hard to reason about.
 Our hypotheses for Sod, Sedov and Bubble experiments prove to be correct.
 However, our intuition turns out to be incorrect for the Cellular experiment.
-Currently, we can apply our tool to C/C++ and Fortran codes, and hence it supports the majority of (large-scale) scientific workloads run on supercomputers.
-While we outlined a few limitations, none of them are fundamental to our approach, and, as we mention, these can be mitigated with additional engineering effort.
-RAPTOR is a step forward towards reasoning about mixed/low precision in our simulations, and we hope that it finds wide adoption in the scientific-computing community.
-We make our RAPTOR tool accessible at https://github.com/RAPTOR-FP/RAPTOR for others to use.
+Currently, we can apply our tool to C/C++ and Fortran codes for CPUs and GPUs, so it supports the majority of (large-scale) scientific workloads run on supercomputers.
+While we outline a few limitations, none of them are fundamental to our approach and we mention how each can be mitigated with additional engineering effort.
+RAPTOR is a step forward towards reasoning about mixed/low precision in scientific simulations, and we hope that it finds wide adoption in the scientific computing community.
+Acknowledgements.
+This work was supported by the Scientific Discovery through Advanced Computing (SciDAC) program via the Office of Nuclear Physics and Office of Advanced Scientific Computing Research in the Office of Science at the U.S.
+Department of Energy.
+This work was supported by Sponsor JST SPRING , Japan Grant Number Grant #JPMJSP2180 and the RIKEN Junior Research Associate Program.
+This work was supported by the National Science Foundation under Grant No.
+2346519, and by the Alan Turing Institute, and the U.S.
+Department of Energy.
+We gratefully acknowledge Hussein Harake and the Swiss National Supercomputing Centre (CSCS) for providing access to computing resources via the RACKlette student cluster, which was instrumental to this work.